# MLLMs-Augmented Visual-Language Representation Learning

### 30 Nov 2023 v1

@@ Abstract
Visual-language pre-training (VLP) have achieved remarkable success in multi-modal tasks, largely attributed to the availability of large-scale image-text datasets. In this work, we demonstrate that multi-modal large language models (MLLMs) can enhance visual-language representation learning by improving data quality. Our approach is simple, utilizing MLLMs to extend multiple captions for each image. To prevent the bias that introduced by MLLMs’ hallucinations and intrinsic caption styles, we propose a “text shearing” to keep the lengths of extended captions identical to the originals. In image-text retrieval, our method consistently obtains 5.6 $\sim$ 35.0% and 16.8 $\sim$ 46.1% improvement on R@1 under the fine-tuning and zero-shot settings, respectively. Notably, our zero-shot results are comparable to fine-tuning on target datasets, which encourages more exploration on the versatile use of MLLMs.
@@ 1 Introduction
Figure 1: The performance in (a) and (b) is zero-shot image-text retrieval results on MSCOCO using BLIP pretrained on CC3M. (c) shows that using a single model to generate captions can easily result in similar text structure and insufficient global considerations.
@@
Visual-language pre-training (VLP) has exhibited remarkable achievements in image-text retrieval[25, 26], image classification[49, 66], vision question answering [3, 31], and image caption generation [32, 66]. The success of VLP can be attributed to the large-scale datasets that are collected from the Internet, such as CC3M [54], CC12M [8], YFCC100M [58], LAION400M [52], etc. However, most of these datasets include a non-negligible portion of noisy and mismatched image-text pairs [25, 60, 29], which largely affects the visual-language representation learning. One of the most straightforward approaches is: utilizing pre-trained models to recognize and remove mismatched pairs based on heuristic rules [17, 1, 7, 38].
@@
These methods indeed reduce the influences of mismatched pairs. However, simply removing mismatched samples leads to a serious problem: the number of training pairs is also reduced. As illustrated in Fig. 0(a), the performance of image-text retrieval consistently drops a lot when the removing ratio is large. Most recently, several works [17, 69, 43] demonstrate that LLM and MLLM can be used as re-writers to improve caption quality without reducing the number of training pairs. Unfortunately, these re-writers inevitably introduce their caption styles, i.e. text structure, which might disrupt the distribution of the original captions and lead to difficulties in learning better visual-language representations.
@@
To investigate the characteristics of previous rewriting works [17, 69, 43], we conduct experiments to evaluate its influences in Fig. 0(b). Experimental results empirically confirm that applying MLLMs to enhance the quality of visual-language datasets is a promising approach, i.e. improving the original performance largely. Nevertheless, excessive re-writing leads to non-trivial performance drops. Therefore, it is valuable to figure out what led to these drops and verify whether there exists bias from MLLMs’ captions.
@@
We explore the underlying reasons for the bias by analyzing the generated captions’ text structures, attention focus, and word frequency. We utilize MiniGPT-4 and LLaVA to caption each image from CC3M, respectively. Here we use the same question {Describe the $\left\langle image\right\rangle$ in English:} to prompt the model and set the same maximum number of generated tokens. As shown in Fig. 0(c) and Fig. 3, we have the following observations. 1). MLLMs indeed have their inherent text structures. 2). Applying a single MLLM makes it hard to provide comprehensive captions. 3). There are great diversities between the word frequency statistics of different MLLMs and raw captions.
@@
Considering the limitations of using a single model to caption the image in terms of diversity, we use multiple MLLMs to enrich visual-language associations from different perspectives, thereby enhancing visual-language representation learning. To mitigate the bias (from text structure and word frequency) introduced by MLLMs, the following specific designs are proposed. First, we reduce the length of captions generated by MLLMs to the same as the raw captions, which is called “text shearing”. It aims to mitigate the influences of intrinsic caption styles and hallucinations from MLLMs. Second, to obtain a comprehensive caption for each image, we keep the raw and extended captions from MLLMs simultaneously.
@@
Our approach exhibits the following characteristics (or advantages): 1). It is compatible with multiple VLP methods like CLIP [49] and BLIP [31], demonstrating significant performance improvements across various downstream tasks. For example, in zero-shot image-text retrieval, our method obtains 16.8$\sim$ 46.1% improvement. 2). For image-text retrieval, without introducing additional training overhead, our zero-shot CLIP outperforms the model fine-tuned on the MSCOCO and Flickr30K respectively by 9.9% and 23.5%. 3). When scaling to large datasets like CC12M and YFCC15M, our method continues to deliver substantial performance improvements.
@@ 2.1 Improving image-text datasets
Numerous studies [18, 42, 20] have emphasized the significance of high-quality image-text datasets in influencing the transfer performance of VLP models in downstream tasks. SemDeDup [1] enhances data efficiency and out-of-distribution performance by identifying and removing semantically duplicated data pairs at the embedding level. Cao et al. [7] introduced an approach to improve dataset quality by removing samples that contain text in images. T-MARS [38] masks the text region in the image and then filters out samples with low CLIP [49] similarity. These methods inevitably lose a lot of visual information when filtering out samples, so some methods try to obtain higher quality data by rewriting captions. Santurkar et al. [51] underscored the importance of synthetic captions and employed pre-trained language models to augment textual content. Gadre et al. [20] introduced the DataComp benchmark for multi-modal datasets. Fan et al. [17] leveraged the in-context learning capacity of large language models to rewrite captions, enriching the language structure while preserving core semantics. Zhu et al. [69] employed ChatGPT [45] and BLIP-2 [32] interactively to generate captions with rich visual information. Nguyen et al. [43] used BLIP-2 [32] to rewrite captions for low-matching image-text pairs and combined original and generated captions for training. Lai et al. [29] proposed enhancing visual information in captions by fusing the original caption with captions generated by LLaVA [36], leveraging large language models.
@@ 2.2 Multimodal large language models
Existing Multimodal Large Language Models (MLLMs) primarily rely on three key technologies: Multimodal Instruction Tuning (M-IT), Multimodal In-Context Learning (M-ICL), and Multimodal Chain-of-Thought (M-CoT) [65]. M-IT facilitates strong transfer performance by fine-tuning the model on datasets with specific instruction formats. Notable models employing this technology include LLaMA-Adapter [67, 21], LLaVA [36, 35], MiniGPT-4 [70], InstructBLIP [13], Qwen-VL [5], and NExT-GPT [62]. M-ICL is a type of analogy learning from a limited number of samples. Models like Flamingo [3], Otter [30], and HuggingGPT [55] have been developed using this approach. M-CoT necessitates models to not only provide answers but also reasoning processes. Representative models include Multimodal-CoT [68] and Visual ChatGPT [61]. These three technologies are not mutually exclusive, and many models effectively combine multiple technologies. Since these MLLMs are often trained on billion-level dataset, they have extremely rich knowledge and excellent visual understanding and expression capabilities, which can be used to caption the images.
Figure 2: The illustration of our method. In terms of images, different MLLMs have different focuses. In terms of language, different MLLMs exhibit distinct idiomatic words and caption styles.
@@ 3.1 Preliminaries
In this section, we briefly introduce the preliminaries of visual-language pre-training. CLIP [49] is a classic work that uses images and text for contrastive learning. It employs an encoder-only architecture, optimizing the encoder through a contrastive loss on image-text pairs. During training, when a batch of $N$ image-text pairs {$x_{I}$ , $x_{T}$ is sampled, the contrastive loss for an image can be defined as follows:
@@
BLIP [31] is an encoder-decoder based visual language pre-training architecture designed to enhance visual language understanding and generation. Its three main pre-training objectives (image-text contrastive loss (ITC), image-text matching loss (ITM), and language modeling loss (LM)) are jointly optimized during the training process. Specifically, ITC is similar to that in CLIP. ITM aims to capture the fine-grained alignment between vision and language. LM is used to optimize the decoder.
As highlighted in many works [66, 41, 33], visual-language pre-training heavily depends on extensive image-text datasets. The presence of low-quality image-text data can significantly compromise the model’s performance. Therefore, improving data quality has become an expected direction.
@@ 3.2 Overview
We demonstrate that different MLLMs can generate accurate and diverse captions in Fig. 2. The goal of this approach is to utilize the MLLMs to improve the data quality in the current image-text datasets. Our approach consists of two processes: multi-view caption extractor and text shearing. Specifically, for each image from an image-text dataset, we first introduce multiple advanced MLLMs to synthesize extended captions. Then, by comprehensively analyzing the characteristics of these extended captions, we have a key observation: MLLMs have their caption styles and might generate hallucinated words. Based on this observation, we propose text shearing to keep the length of extended captions the same as the raw one. After these processes, the raw and extended captions with corresponding images are used for standard visual-language pertaining methods.
@@ 3.3 Multi-View caption extractor
How to use these extended captions? After obtaining these extended captions, one of the most straightforward ideas is to replace the raw texts with a set of new captions for training the visual-language model. Its effectiveness has been evaluated in many previous works [69, 43]. However, we have a concern about the operation in terms of the difference between raw and new captions. To investigate this, we count the frequency of the most occurring nouns of the extended captions (extracted from images in CC3M by MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, examples are shown in Fig. 2) and present the results in Fig. 3. One can find that different MLLMs have different common words in the generated captions.
@@
Figure 4: Average length of captions generated using different MLLMs.
Based on the above analysis, we plan to use all the extended captions for training to improve the diversity of captions. However, there is another concern about the difference in the length of the raw and extended captions. We present the average length comparisons of these captions in Fig. 4. These captions are generated by MLLMs under default settings.It can be seen that the generated captions by MiniGPT-4 and LLaVA-1.5 are notably longer, i.e. 90 tokens. While raw caption is shorter, i.e. 15 tokens. To investigate the influence of varying caption lengths on pre-training, we visualize the longer caption generated by MiniGPT-4 in Figure 5. We find that longer captions indeed produce hallucinations that do not match the image, and these hallucinations often appear at the end of the caption, which has an impact on representation learning.
@@ 3.4 Text Shearing
In order to eliminate the impact of caption length, we propose text shearing to make the length of the extended captions as same as the raw ones. Specifically, we add a token number $T$ as a limit when using MLLMs to generate captions. In order to prevent the generated caption from having incomplete expressions, we intercept the first complete clause in the generated caption. Therefore, Eq. 2 can be rewritten as:
@@
For the setting of $T$, we use a simple but effective strategy. We set it to the average length of the original captions to mimic human annotation. By controlling the length of generated captions, we hope to effectively reduce the impact of MLLM’s hallucination and its intrinsic caption style.
Then, we jointly use original captions and newly generated captions to construct a new dataset. The new image-text pairs can be expressed as $\{x_{I},x_{T}^{\prime}\}=\{x_{I},c_{1}^{i^{\prime}}\},\{x_{I},c_{2}^{i^{\prime}}\}\cdots,\{x_{I},c_{k}^{i^{\prime}}\}$, where $k$ represents the number of MLLM. We use the enhanced dataset for subsequent standard visual-language pre-training. Taking CLIP as an example, the loss on the image can be written as:
@@
Table 2: The zero-shot evaluation results of our method on 15 common classification datasets and ImageNet with CLIP pre-trained on CC3M and CC12M. The best results are highlighted in bold. On CC3M, our method demonstrates an average improvement of 13.4 on 15 datasets, and 13.1 on ImageNet. The different results on CLIP are due to the different training data downloaded from the Internet.

### 13 Mar 2024 v3

@@ Abstract
Visual-language pre-training has achieved remarkable success in many multi-modal tasks, largely attributed to the availability of large-scale image-text datasets. In this work, we demonstrate that Multi-modal Large Language Models (MLLMs) can enhance visual-language representation learning by establishing richer image-text associations for image-text datasets. Our approach is simple, utilizing MLLMs to extend multiple diverse captions for each image. To prevent the bias introduced by MLLMs’ hallucinations and monotonous language styles, we propose “text shearing" to maintain the quality and availability of extended captions. In image-text retrieval, without introducing additional training cost, our method consistently obtains 5.6 $\sim$ 35.0% and 16.8 $\sim$ 46.1% improvement on Recall@1 under the fine-tuning and zero-shot settings, respectively. Notably, we obtain zero-shot results that are comparable to fine-tuning on target datasets, which encourages more exploration of the versatile use of MLLMs. The datasets and codes are available at: https://github.com/lyq312318224/MLLMs-Augmented.
Keywords:
Visual-language pre-training Representation learning
**footnotetext: equal contributionfootnotetext: project leadfootnotetext: corresponding author
@@ 1 Introduction
Figure 1: The performance in (a) and (b) is zero-shot image-text retrieval results on MSCOCO using pretrained BLIP. (c) shows that using a single model to generate captions can easily result in similar text structure and insufficient global considerations.
@@
Visual-language pre-training has achieved remarkable success in image-text retrieval[25, 26], image classification[49, 66], vision question answering [3, 32], and image caption generation [31, 66]. This success can be attributed to the large-scale datasets collected from the Internet, such as CC3M [54], CC12M [8], YFCC100M [58], LAION400M [52], etc. However, most of these datasets include a non-negligible portion of noisy and mismatched image-text pairs [25, 60, 29], which largely affects the visual-language representation learning. One of the most straightforward approaches is: utilizing pre-trained models to recognize and remove mismatched pairs based on heuristic rules [17, 1, 7, 38].
@@
These methods indeed reduce the influence of mismatched pairs. However, simply removing mismatched pairs leads to a serious problem: the number of training pairs is also reduced. As illustrated in Figure 0(b), the performance of image-text retrieval consistently drops a lot when the removal ratio is large. Most recently, several works [17, 70, 42] demonstrate that LLM and MLLM can be used as re-writers to improve caption quality without reducing the number of training pairs. Unfortunately, these re-writers inevitably introduce their caption styles, i.e. text structure, which might disrupt the distribution of the original captions and lead to difficulties in learning better visual-language representations.
@@
To investigate the characteristics of previous rewriting works [17, 70, 42], we conduct experiments to evaluate their influences in Figure 0(b). Experimental results empirically confirm that applying MLLMs to enhance the quality of visual-language datasets is a promising approach, i.e. improving the original performance largely. Nevertheless, excessive re-writing leads to non-trivial performance drops. Therefore, it is valuable to figure out what led to these drops and verify whether there exists bias from synthetic captions.
@@
We explore the underlying reasons for the bias by analyzing the generated captions’ text structures, attention focus, and word frequency. We utilize MiniGPT-4 and LLaVA to caption each image from CC3M, respectively. Here we use the same question {Describe the $\left\langle image\right\rangle$ in English:} to prompt the model and set the same maximum number of generated tokens. As shown in Figure 0(c) and Figure 3, we have the following observations. 1). MLLMs indeed have their inherent text structures and focus. 2). Applying a single MLLM makes it hard to provide comprehensive captions. 3). There are great diversities between the word frequency statistics of different MLLMs and raw captions.
@@
Considering the limitations of a single model in capturing diverse image captions, we employ multiple MLLMs to enrich visual-language associations from different perspectives, thereby improving visual-language representation learning during pre-training. To improve the quality and availability of synthetic captions, the following specific designs are proposed. First, we crop the captions generated by MLLMs to match the length of the original captions, which is called “text shearing”. This not only reduces the repetitive occurrence of common words in synthetic captions, alleviating the caption collapse problem identified in [60], but also preserves the semantic concepts closest to the image, thereby mitigating the impact of hallucinations typically appearing in the later part of the generated text, as verified in [69]. Second, to obtain a set of comprehensive captions for each image, we keep the raw and extended captions from MLLMs simultaneously for standard visual-language pre-training.
@@
Our approach exhibits the following characteristics: 1). It is compatible with multiple visual-language pre-training frameworks like CLIP [49] and BLIP [32], demonstrating significant performance improvements across various downstream tasks without introducing additional training overhead. For example, in MSCOCO and Flickr30K’s zero-shot image-text retrieval, our method obtains 16.8$\sim$ 46.1% Recall@1 improvement. In zero-shot image classification, our method achieves an average performance improvement of 13.4 on 15 common classification datasets and 13.1 on ImageNet [14]. 2). For image-text retrieval, our zero-shot CLIP outperforms the vanilla CLIP fine-tuned on the MSCOCO and Flickr30K respectively by 9.9% and 23.5%. (The above results are all based on pre-training on CC3M.) 3). When scaling to large datasets like CC12M and YFCC15M, our method continues to deliver substantial performance improvements.
@@ 2.1 Improving image-text datasets
Numerous studies [18, 43, 20] have emphasized the significance of high-quality image-text datasets in influencing the transfer performance of visual-language pre-training in downstream tasks. SemDeDup [1] enhances data efficiency and out-of-distribution performance by identifying and removing semantically duplicated data pairs at the embedding level. Cao et al. [7] introduced an approach to improve dataset quality by removing samples that contain text in images. T-MARS [38] masks the text region in the image and then filters out samples with low CLIP [49] similarity. These methods inevitably lose a lot of visual information when filtering out samples, so some methods try to obtain higher-quality data by rewriting captions. Santurkar et al. [51] underscored the importance of synthetic captions and employed pre-trained language models to augment textual content. Gadre et al. [20] introduced the DataComp benchmark for multi-modal datasets. Fan et al. [17] leveraged the in-context learning capacity of large language models to rewrite captions, enriching the language structure while preserving core semantics. Zhu et al. [70] employed ChatGPT [45] and BLIP-2 [31] interactively to generate captions with rich visual information. Nguyen et al. [42] used BLIP-2 [31] to rewrite captions for low-matching image-text pairs and combined original and generated captions for training. Lai et al. [29] proposed enhancing visual information in captions by fusing the original caption with captions generated by LLaVA [36], leveraging large language models.
@@ 2.2 Multimodal large language models
Existing Multimodal Large Language Models (MLLMs) primarily rely on three key technologies: Multimodal Instruction Tuning (M-IT), Multimodal In-Context Learning (M-ICL), and Multimodal Chain-of-Thought (M-CoT) [65]. M-IT facilitates strong transfer performance by fine-tuning the model on datasets with specific instruction formats. Notable models employing this technology include LLaMA-Adapter [67, 21], LLaVA [36, 35], MiniGPT-4 [71], InstructBLIP [13], Qwen-VL [5], and NExT-GPT [62]. M-ICL is a type of analogy learning from a limited number of samples. Models like Flamingo [3], Otter [30], and HuggingGPT [55] are developed using this approach. M-CoT necessitates models to not only provide answers but also reasoning processes. Representative models include Multimodal-CoT [68] and Visual ChatGPT [61]. These three technologies are not mutually exclusive, and many models effectively combine multiple technologies. Since these MLLMs are often trained on the billion-level dataset, they have extremely rich knowledge and excellent visual understanding and expression capabilities, which can be used to caption the images.
Figure 2: The illustration of our method. Different MLLMs jointly construct one-to-many image-text pairs with richer semantic associations.
@@ 3.1 Preliminaries
In this section, we briefly introduce the preliminaries of visual-language pre-training. CLIP [49] is a classic work that uses image-text pairs for contrastive learning. It employs an encoder-only architecture, optimizing the encoder through a contrastive loss on image-text pairs. During training, when a batch of $N$ image-text pairs {$x_{I}$ , $x_{T}$ is sampled, the contrastive loss for an image can be defined as follows:
@@
BLIP [32] is an encoder-decoder based visual-language pre-training architecture. Its three main pre-training objectives (image-text contrastive loss (ITC), image-text matching loss (ITM), and language modeling loss (LM)) are jointly optimized during the training process. Specifically, ITC is similar to that in CLIP. ITM aims to capture the fine-grained alignment between vision and language. LM is used to optimize the decoder.
As highlighted in many works [66, 41, 33], visual-language pre-training heavily depends on extensive image-text datasets. The presence of low-quality image-text data can significantly compromise the model’s performance. Therefore, improving the multi-modal dataset has become an expected direction.
@@ 3.2 Overview
In this work, we employ multiple advanced MLLMs to augment the visual-language representation learning from a data-centric perspective. We demonstrate that different MLLMs can generate accurate and diverse captions in Figure 2. The goal of this approach is to utilize the MLLMs to establish richer image-text associations in the current datasets. Our approach consists of two processes: multi-view caption extractor and text shearing. Specifically, for each image from an image-text dataset, we first introduce multiple advanced MLLMs to synthesize extended captions. Then, by comprehensively analyzing the characteristics of these extended captions, we have a key observation: MLLMs have their caption styles and might generate hallucinated content. Based on this observation, we propose text shearing to maintain the quality and availability of extended captions. After these processes, the raw and extended captions with corresponding images are jointly used for standard visual-language pre-training.
@@ 3.3 Multi-View caption extractor
How to use these extended captions? After obtaining these extended captions, one of the most straightforward ideas is to replace a set of raw captions with new captions for training. Its effectiveness has been evaluated in many previous works [70, 42]. However, we have a concern about the operation in terms of the difference between raw and new captions. To investigate this, we count the frequency of the most occurring nouns of the extended captions (extracted from images in CC3M by MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, examples are shown in Figure 2) and present the results in Figure 3. One can easily find that different MLLMs output different common words in the generated captions. These diverse words can greatly expand the linguistic concepts of the dataset.
@@
(a) Average length of captions generated using different MLLMs.
@@
Figure 4: Analysis of extended captions.
Based on the above analysis, we plan to use all the extended captions for training to improve the diversity of captions. However, there is another concern about the difference in the length of the raw and extended captions. We present the average length comparisons of these captions in Figure 3(a). These captions are generated by MLLMs under default settings.It can be seen that the generated captions by MiniGPT-4 and LLaVA-1.5 are notably longer, i.e. 90 tokens. While raw caption is shorter, i.e. 15 tokens. To investigate the influence of varying caption lengths on pre-training, we visualize the longer caption generated by MiniGPT-4 in Figure 3(b). We find that longer captions indeed produce hallucinations that do not match the image, and these hallucinations often appear in the later part of the generated text, which has an impact on representation learning. The same conclusion is also verified in  [69]. It states that as generation progresses, the accumulation of past hallucinatory information and uncertainty can derail the model and lead to new hallucinations.
@@ 3.4 Text Shearing
To eliminate the impact of excessive caption length, we propose text shearing to make the length of the extended captions as same as the raw ones. Specifically, we add a token number $T$ as a limit when using MLLMs to generate captions. To prevent the generated captions from having incomplete expressions, we intercept the first complete clause in the generated caption. Therefore, Eq. 2 can be rewritten as:
@@
For the setting of $T$, we use a simple but effective strategy. We set it to the average length of the original captions to mimic human annotation. By controlling the length of generated captions, we hope to effectively reduce the impact of MLLMs’ hallucinations and their monotonous language styles.
Then, we jointly use original captions and newly generated captions to construct a new dataset. The new image-text pairs can be expressed as $\{x_{I},x_{T}^{\prime}\}=\{x_{I},c_{1}^{i^{\prime}}\},\{x_{I},c_{2}^{i^{\prime}}\}\cdots,\{x_{I},c_{k}^{i^{\prime}}\}$, where $k$ represents the number of MLLMs. We use the enhanced dataset for subsequent standard visual-language pre-training. Taking CLIP as an example, the loss on the image can be written as:
@@
4 Experiments
4.1 Pre-training Details
Our method is implemented in Pytorch [47] and trained on a node equipped with 8 NVIDIA A100 GPUs. For visual language pre-training, we follow the implementation of CLIP [49] and BLIP [32] respectively. For CLIP, we train the model based on the open source code OpenCLIP [24]. During training, our batch size is set to 320. The number of epochs is set to 6 to better align the training overhead. The model architecture is set as VIT-B-16. Other parameters are the same as the default values in OpenCLIP [24]. For BLIP, we use a ViT-B-16 [16, 59] pre-trained on ImageNet [14] as the image encoder and $\text{Bert}_{\text{base}}$ [15] as the text encoder. Regarding image preprocessing, we employ random cropping to achieve $224\times 224$ during pre-training and increase the image resolution to $384\times 384$ during fine-tuning. We apply RandomAugment [12] for image augmentation. For optimization, we utilize an AdamW [37] optimizer with a learning rate warm-up to 3e-4, combined with linear decay at a rate of 0.9. Our training batch size is set to 1280, and the number of training epochs is 4.
Regarding the usage of multi-modal large language models, we employed Mini-GPT4-Vicuna13B [71], Otter-Image-MPT7B  [30], Qwen-VL-Chat [5], and LLaVA-v1.5-13B [35]. During caption generation, we impose a maximum token limit of 30 to control the caption length. For the synthetic caption generation, we use beam search with the number of beams equal to 1.
We use CC3M [54], CC12M [8] and YFCC15M as pre-training datasets. Among them, YFCC15M is a subset of YFCC100M [58] that contains English descriptions of images. Due to the inaccessibility of some images, the version of the dataset we used contains fewer images than the original dataset. Among them, on CC3M, compared with the 3.3M version of the original dataset, the version we used contains 2.6M image-text pairs; on CC12M, compared with the 12.4M version of the original dataset, the version we used contains 11.1M image-text pairs; On YFCC15M, compared to the 15M of the original dataset, our version contains 14.8M image-text pairs.
@@ 4.1 Pre-training Details
4.2 Evaluation
We evaluate the performance of the multi-modal task in Table 1 and Table 4. We also provide results on the image classification task in Table 2 and Table 3.
Image-Text Retrieval. Image-text retrieval serves as a crucial metric for evaluating the bridging capability between different modalities in a model. We pre-train CLIP and BLIP on CC3M, conducting image-to-text retrieval (TR) and text-to-image retrieval (IR) testing on the MSCOCO [34] and Flickr30K [48] datasets. Results in Table 1 demonstrate significant performance improvements in both zero-shot and fine-tuning retrievals for both CLIP and BLIP. Notably, our approach enables CLIP’s zero-shot retrieval performance to outperform models finetuned on the target dataset. Specifically, when using CLIP for zero-shot retrieval on MSCOCO, the R@1 of TR and IR increase by 27.2 and 19.4, respectively. For zero-shot retrieval on Flickr30K, the R@1 of TR and IR also improves by 46.1 and 35.4, respectively. Similarly, our method also achieves an improvement of 16.8$\sim$23.4 when using BLIP for zero-shot retrieval. These significant performance improvements indicate that the pre-trained model learns better image-text representation.
Table 2: The zero-shot evaluation results of our method on 15 common classification datasets and ImageNet with CLIP pre-trained on CC3M and CC12M. The best results are highlighted in bold. On CC3M, our method demonstrates an average improvement of 13.4 on 15 datasets and 13.1 on ImageNet. The different results on vanilla CLIP are due to the different training data downloaded from the Internet.